OpenStudy (anonymous):

I need to find the limit of the Function ((1/3+x)-(1/3))/x as x aproches 0

OpenStudy (anonymous):

what are you allowed to use?

OpenStudy (anonymous):

\[\frac{\frac{1}{3}+x-\frac{1}{3}}{x}\] like this?

myininaya (myininaya):

\[\frac{\frac{1}{3+x}-\frac{1}{3}}{x}?\]

OpenStudy (anonymous):

ooh, that makes more sense

myininaya (myininaya):

lol

OpenStudy (anonymous):

go ahead. and when you are done, tell me where i can find those nice fan testimonials

OpenStudy (anonymous):

yes Myininaya thats it

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\frac{1}{3+x}-\frac{1}{3}}{x} \cdot \frac{(3+x)(3)}{(3+x)(3)}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{3-(3+x)}{3x(3+x)}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{-x}{3x(3+x)}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{- \not{x}}{3 \not {x} (3+x)}=\lim_{x \rightarrow 0}\frac{-1}{3(3+x)}\]

myininaya (myininaya):

now plug in zero

myininaya (myininaya):

\[\frac{-1}{3(3+0)}=\frac{-1}{9}\]

OpenStudy (anonymous):

Thank you so much and sorry about the equation confusion I just started using the site

myininaya (myininaya):

satellite you want to see your fan testimonials or do you want to leave others a fan testimonial? just go to your profile or someone's to read theirs or to leave one

myininaya (myininaya):

its cool vinder

myininaya (myininaya):

did you follow everything i did vinder?

OpenStudy (anonymous):

oh so now you want to clear the fractions. before you did the arithmetic. i will do it that way \[\frac{\frac{3-(3+x)}{(3+x)3}}{x}\] \[\frac{-x}{(3+x)3x}\] \[\frac{-1}{(3+x)3}\] and \[\lim_{x\rightarrow }\frac{-1}{(x+3)3}=\frac{-1}{3\times 3}=-\frac{1}{9}\]

OpenStudy (anonymous):



myininaya (myininaya):

satellite combine the top fractions first which is fine too

OpenStudy (anonymous):

that is what you used to do i remember. i want to see them and leave them both!

OpenStudy (anonymous):

Yeah I forgot how to simplify multiple fractions in the numerator thank you so much

myininaya (myininaya):

yes i used to do it that way but i like this way better

OpenStudy (anonymous):

i like it better for clearing complex fractions but for some reason here i like the artithmetic